OpenStudy (cobaltthunder267):

Find an exact value of \[\cos \frac{ \pi }{ 12 }\]

OpenStudy (javalos01):

need hellp

OpenStudy (javalos01):

it will help u understand

OpenStudy (anonymous):

Let \(x=\cos\dfrac{\pi}{12}\). Then \(x^2=\cos^2\dfrac{\pi}{12}\). Using the half angle identity, \[\cos^2\theta=\frac{1+\cos2\theta}{2}\] we have \(x^2=\dfrac{1+\cos\dfrac{\pi}{6}}{2}\). Find the value of \(\cos\dfrac{\pi}{6}\) and take the square root to find \(x\).

OpenStudy (cobaltthunder267):

@javalos01 Ok! That did help a lot! Thank you :) There's a whole bunch of results for this on Google, but they're all different so I wanted to confirm. @SithsAndGiggles Thanks!! That helped a bunch, too. I think I've got it now. Thank you both!!

OpenStudy (javalos01):

i know thats a good one

OpenStudy (javalos01):

hey ill fan u

OpenStudy (cobaltthunder267):

Lol ok. Right back at 'cha ;)